# Precision Determination of Scintillation Screen Parameters from Annual Modulation Measurement of Pulsar Scintillation Arc Curvature with the FAST Telescope

### v1

@@ 3.2 Dynamic Spectra
This second masking process removed only about 2–3% of the dynamic spectrum data, but significantly improved the quality of the secondary spectrum for some observations.
3.3 Secondary Spectra
We obtained the secondary spectra by performing a fast Fourier transform (FFT) on the dynamic spectra and taking the square of the modulus of the results. Before performing the FFT, we interpolated each dynamic spectrum onto a grid of equal wavelength spacing, in order to eliminate the dependence of scintillation arc curvature on observing frequency (e.g., Reardon et al. 2020; Ocker et al. 2024). The frequency-independent arc curvature $\eta$, such that $f_{\lambda}$ = $\eta f_{t}^{2}$ (where $f_{\lambda}$ is the differential time delay of wavelength conjugation), is given by
$\eta=\frac{\text{d}_{\text{psr}}s(1-s)}{2\text{V}_{\text{sr}}^{2}}.$ (10)
@@ 3.3 Secondary Spectra
$i(u)=\begin{cases}u&(u\leq m)\\ 2m-u&(u>m)\end{cases}\ \ ,$ (12)
$\text{and}\quad j(v)=\begin{cases}v&(v\leq n)\\ 2n-v&(v>n)\end{cases}\ \ .$
This approach preserves more signal at the cost of mixing the parities of the signal. As a consequence, our secondary spectra are symmetric around $f_{t}=0$ and unsuitable for identifying substructures on the parabolic arcs.
3.4 Scintillation Arc Search and Curvature Measurement
We rescaled the secondary spectrum intensity on a logarithmic scale, using dB as the unit of intensity. We set the noise floor to 0 dB, and any data below the noise floor were excluded from the scintillation arc search and curvature calculation.
To search for the scintillation arc, we computed the mean intensity of the secondary spectrum along parabolic arc paths. For low arc curvatures (below 100 $\mathrm{m}^{-1}\ \mathrm{mHz}^{-2}$), the mean intensity was calculated out to $f_{\lambda}$ = 5000 $\mathrm{m}^{-1}$. We identified the peak of the resulting mean-intensity curve as the signal associated with the scintillation arc. This process is known as the generalized Hough transform (Ballard, 1981; Bhat et al., 2016). The generalized Hough transform and its variation (e.g., Xu et al. 2018; Reardon et al. 2020) are common methods for searching and measuring scintillation arcs. Before identifying the arc signal, we reduce noise using a low-pass filter.
@@ 3.4 Scintillation Arc Search and Curvature Measurement
A 84.8$\pm$4.4 - 78.7$\pm$1.8 78.9$\pm$3.1 -
B 122.9$\pm$4.1 123.0$\pm$7.4 125.4$\pm$4.9 - -
C 221.1$\pm$12.5 221.4$\pm$9.3 231.8$\pm$14.3 241.1$\pm$11.0 -
D 635$\pm$23 607$\pm$21 710$\pm$39 901$\pm$27 -
E 1871$\pm$78 1938$\pm$71 - 6697$\pm$256 4319$\pm$287
F (?) 7224$\pm$872 14893$\pm$1719 2141$\pm$79 - -
Note. — For arc group F, due to limited observation data and significant curvature variations, we cannot confirm whether these arcs originate from the same screen.
Figure 2: Ten scintillation arcs were observed in pulsar B1237+25. Among them, nine scintillation arcs (excluding arc I) were identified during the observation on January 18, 2020. The left panel shows the secondary spectrum of that day, while the right panel shows the fitted scintillation arcs, labeled from A to J (excluding I) in order of increasing curvature.
@@ 4 Results
We observed scintillation arcs in all observations (as shown in Table 3, 4, and 5). The observed scintillation arcs varied in structure, with some being more diffuse (e.g., PSR B1842+14, as shown in Figure 4) and others being finer (e.g., the scintillation arc G of PSR B1237+25, as shown in Figure 2 and 3). We note that except for arc F of pulsar B2021+51, all other scintillation arcs exhibited sharp parabolic arc characteristics, consistent with strongly anisotropic scintillation screens. For each anisotropic arc, we modeled its annual modulation using the Markov Chain Monte Carlo (MCMC) method to calculate the fractional distance s of the scintillation screen, the velocity component of the scintillation screen along the scattering direction $\text{V}_{{\text{screen},~{}\zeta}}$, and the scattering direction $\mathbf{n_{\text{screen}}}$, as shown in Table 6. We express the measured direction of the scattering direction using the azimuth angle $\zeta$ as degrees east of north. The MCMC sampling was performed using the Python package emcee (Foreman-Mackey et al., 2013), which efficiently explores the parameter space to generate posterior distributions for the parameters.
@@
However, we were unable to calculate all three parameters for every scintillation screen when modeling the modulation. As discussed in Section 2, in cases of anisotropic scattering, the direction of the scattering direction of the scintillation screen is the most easily constrained parameter, as it only requires knowledge of the point of minimum curvature. The fractional distance s is the next most easily constrained, while the screen’s projected velocity $\text{V}_{{\text{screen},~{}\zeta}}$ is the most difficult to constrain. For some screens, we could only calculate partial parameters based on certain assumptions. Below are the situations we encountered and our approaches, respectively:
@@
1. Substantial errors. The relative errors in the curvature measurements of the scintillation arc H for PSR B1237+25 are substantial, resulting in poor constraints on the screen parameters. In this case, considering that the projected velocity $\text{V}_{{\text{screen},~{}\zeta}}$ of the screen is often relatively small and is the most difficult to constrain, we assumed its value to be 0 and, after modeling, could only obtain estimates of s and $\zeta$.
@@
2. Weak modulation. When the scintillation screen is close to the pulsar, i.e., $\text{s}{\text{V}_{\text{Earth}}}\ll(1-\text{s})\text{V}_{\text{psr}}$, the modulation amplitude of the arc curvature is not significantly greater than the measurement error. For example, the difference between the maximum and minimum curvature of scintillation arc A for PSR B1237+25 is only 0.15 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$, while the measurement error ranges from 0.05 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$ to 0.13 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$ (see Table 3). Scintillation arcs D and F of PSR B1237+25 (see Table 3) and scintillation arcs A, B, and C of PSR B2021+51 (see Table 5) also exhibit this characteristic. In such cases with small modulation amplitude, we could only constrain one screen parameter. Considering that $\text{V}_{{\text{screen},~{}\zeta}}$ is often relatively small, and these scintillation screens are relatively close to the pulsar, satisfying $\text{V}_{{\text{screen},~{}\zeta}}\ll(1-\text{s})\text{V}_{\text{psr}}$, we assumed $\text{V}_{{\text{screen},~{}\zeta}}$ to be 0. We also assumed the main scattering direction $\mathbf{n_{\text{screen}}}$ is aligned with the effective velocity (see e.g. Sprenger et al. 2020).
$\mathbf{V}_{\mathrm{eff}}\equiv-\frac{1}{\mathrm{s}}\mathbf{V}_{\mathrm{screen}}+\mathbf{V}_{\mathrm{Earth}}+\frac{1-\mathrm{s}}{\mathrm{s}}\mathbf{V}_{\mathrm{psr}}.$ (13)
@@
3. Insufficient number of detections. Some scintillation arcs could only be identified in observations conducted on dates with close day-of-year positions. Scintillation arcs B and E of PSR B1237+25 (see Table 3) were only observed on January 17, 2020, and January 14, 2023, which is effectively equivalent to having only a single observation. In such cases, we assumed $\text{V}_{{\text{screen},~{}\zeta}}$ to be 0 and $\mathbf{n_{\text{screen}}}$ to be aligned with $\mathbf{V_{\text{eff}}}$ to calculate the value of s. We obtained two solutions for s: the maximum value of s if the scintillation screen is close to the pulsar, and the minimum value of s if the scintillation screen is close to Earth.
4. Insufficient sampling of the modulation curve. For scintillation arc I of PSR B1237+25, two of the four observations were conducted on dates with close day-of-year positions (November 21, 2019, and November 16, 2022), lacking sufficient phase information about the annual modulation of curvature. This made constraining the screen parameters difficult. In this case, we assumed $\text{V}_{{\text{screen},~{}\zeta}}$ to be 0, and obtained estimates of s and $\zeta$. For PSR B1237+25, the scintillation arc C was observed twice with about a 129-day interval in day-of-year positions, and their measured curvature values were very close, even less than the measurement error. Arc C could be a weakly modulated arc, or it might be a strongly modulated arc that happened to be observed on two dates with very similar curvature values. Therefore, we assumed $\text{V}_{{\text{screen},~{}\zeta}}$ to be 0 and $\mathbf{n_{\text{screen}}}$ to be aligned with $\mathbf{V_{\text{eff}}}$ to calculate the value of s. We obtained two solutions for s. The scintillation arc F of PSR B2021+51 had fewer observation data points and large differences in measured curvature. We could not determine whether this group of scintillation arcs originated from the same screen, so we did not fit it.
@@
Table 6: Parameters of scintillation screens observed in the directions of pulsars B1237+25, B1842+14, and B2021+51. The parameter s represents the fractional distance of the scintillation screen, $\text{d}_{\text{screen}}$ represents the distance from the scintillation screen to the Earth, $\zeta$ represents the azimuth angle (degrees east of north) of the scattering direction, and $\text{V}_{{\text{screen},~{}\zeta}}$ represents the velocity component along the scattering direction. For screens without a $\zeta$ value, we assume their scattering direction and effective velocity are aligned. For screens without a $\text{V}_{{\text{screen},~{}\zeta}}$ value, we assume that their $\text{V}_{{\text{screen},~{}\zeta}}$ is zero.
Pulsar Screen ID s $\text{d}_{\text{screen}}$ (pc) $\zeta$ () $\text{V}_{{\text{screen},~{}\zeta}}$ (km/s)
PSR B1237+25 A $0.0435_{-0.0027}^{+0.0031}$ $824^{+61}_{-53}$
@@
F $0.512^{+0.018}_{-0.017}$ $421^{+35}_{-31}$
G $0.691^{+0.026}_{-0.028}$ $267^{+32}_{-28}$ $124.7^{+5.1}_{-5.2}$ $-1.2^{+8.6}_{-9.5}$
H $0.632^{+0.096}_{-0.151}$ $321^{+130}_{-88}$ $171.1^{+10.0}_{-10.5}$
I $0.756^{+0.114}_{-0.224}$ $214^{+191}_{-102}$ $175.1^{+11.6}_{-12.7}$
J $0.723^{+0.069}_{-0.099}$ $241^{+86}_{-63}$ $63.7^{+8.4}_{-7.3}$ $-41^{+22}_{-37}$
PSR B1842+14 - $0.85^{+0.05}_{-0.10}$ $2.4^{+2.1}_{-1.2}$$\times 10^{2}$ $122^{+11}_{-12}$ $19^{+11}_{-11}$
PSR B2021+51 A $0.0368^{+0.0062}_{-0.0047}$ $1.93^{+0.31}_{-0.24}$$\times 10^{3}$
B $0.0567^{+0.0093}_{-0.0072}$ $1.88^{+0.31}_{-0.23}$$\times 10^{3}$
C $0.100^{+0.015}_{-0.012}$ $1.80^{+0.29}_{-0.22}$$\times 10^{3}$
@@
E $0.556^{+0.041}_{-0.038}$ $887^{+167}_{-132}$ $20.3^{+4.3}_{-4.3}$ $-12.1^{+3.7}_{-3.8}$
Figure 6: Screen distance estimates for B1237+25 (top), B1842+14 (middle), and B2021+51 (bottom). The scintillation screen labeled “No assumption” indicates their positions were precisely calculated with all three parameters determined. The scintillation screen labeled “One assumption” indicates their positions were calculated based on one assumption that the velocity is zero. The scintillation screen labeled “Two assumptions” indicates their positions were calculated based on two assumptions: zero velocity and scattering direction aligned with the effective velocity. For pulsars with multiple scintillation screens, each screen has been marked with its ID. For PSR B1237+25, scintillation screens B, C, and E have two sets of distance solutions. For solutions close to the Sun, we label them as $\text{B}_{1}$, $\text{C}_{1}$, and $\text{E}_{1}$, while for solutions close to the pulsar, we label them as $\text{B}_{2}$, $\text{C}_{2}$, and $\text{E}_{2}$. The black dashed line represents the distance of the Local Bubble inner surface (Pelgrims et al., 2020; Pelgrims, 2022).
@@
4.1 PSR B1237+25
We identified 10 scintillation arcs in the direction of pulsar B1237+25, as shown in Table 3. Among them, 9 scintillation arcs could be identified in the observation data from January 18, 2020, as shown in Figure 2.
Scintillation arc G had the highest brightness and was identified in all 10 observations. We measured and modeled the curvature modulation of arc G (as shown in Figure 8). We calculated that the fractional distance s of scintillation screen G is $0.691^{+0.026}_{-0.028}$ (corresponding to a distance of $267^{+32}_{-28}$ pc from Earth), the scattering direction $\zeta$ is $124.7^{+5.1}_{-5.2}$ , and the velocity of the scintillation screen along the scattering direction $\text{V}_{{\text{screen},~{}\zeta}}$ is $-1.2^{+8.6}_{-9.5}$ km/s. We also modeled and measured the curvature modulation of scintillation arcs J and I, successfully calculating all three parameters.
The measurement error of the curvature of scintillation arc H was relatively large, making it difficult to constrain the scintillation screen parameters, especially the $\text{V}_{{\text{screen},~{}\zeta}}$. Therefore, we assumed $\text{V}_{{\text{screen},~{}\zeta}}$ to be 0 and only calculated the fractional distance s and the scattering direction. Scintillation arc I was identified in only 4 observations, with two occurring on dates in close proximity, preventing us from obtaining sufficient phase information about the curvature modulation. Hence, we also assumed $\text{V}_{{\text{screen},~{}\zeta}}$ to be 0.

### v2

@@ III.2 Dynamic Spectra
This second masking process removed only about 2–3% of the dynamic spectrum data, but significantly improved the quality of the secondary spectrum for some observations. An example comparing the results before and after this process is shown in Figure 8.
III.3 Secondary Spectra
We obtained the secondary spectra by performing a fast Fourier transform (FFT) on the dynamic spectra and taking the square of the modulus of the results. Before performing the FFT, we interpolated each dynamic spectrum onto a grid of equal wavelength spacing, in order to eliminate the dependence of scintillation arc curvature on observing frequency (e.g., Fallows et al. 2014; Reardon et al. 2020; Ocker et al. 2024). Note that this method smears discrete features along the arc, making it unsuitable for studies of discrete features such as inverted arclets (Sprenger et al., 2020). However, this does not hinder the detection and fitting of arcs. The frequency-independent arc curvature $\eta$, such that $f_{\lambda}$ = $\eta f_{t}^{2}$ (where $f_{\lambda}$ is the differential time delay of wavelength conjugation), is given by
$\eta=\frac{d_{\text{psr}}s(1-s)}{2v_{\text{sr}}^{2}}.$ (10)
@@ III.3 Secondary Spectra
[…]
This approach retains more of the signal at the cost of mixing signals of different parities. The secondary spectra obtained through this method are symmetric about $f_{t}=0$ and cannot preserve asymmetries in the data, potentially obscuring some arcs when multiple fine arcs are present. Through careful comparison of secondary spectra obtained with and without this method, we find that none of the arcs in our study are obscured. Note that this method is not suitable for fitting parabolic arcs with apexes offset from the origin or for identifying arc substructures. The parabolic arcs in our study show no detectable apex offset (see APPENDIX A for detailed discussion). This method reduces the background noise in the secondary spectra, as shown in Figure 8.
III.4 Scintillation Arc Search and Curvature Measurement
We rescaled the secondary spectrum intensity on a logarithmic scale, using dB as the unit of intensity. We set the noise floor to 0 dB, and any data below the noise floor was excluded from the scintillation arc search and curvature calculation.
To search for scintillation arcs, we computed the mean intensity of the secondary spectrum along parabolic arc paths. For low arc curvatures (below 100 $\mathrm{m}^{-1}\ \mathrm{mHz}^{-2}$), the mean intensity was calculated out to $f_{\lambda}$ = 5000 $\mathrm{m}^{-1}$. We identified the peak of the resulting mean-intensity curve as the signal associated with the scintillation arc. This process is known as the generalized Hough transform (Ballard, 1981; Bhat et al., 2016). The generalized Hough transform and its variation (e.g., Xu et al. 2018; Reardon et al. 2020) are common methods for searching and measuring scintillation arcs. Before identifying the arc signal, we reduce noise using an FFT-based low-pass filter to smooth the intensity curve along each parabolic path. The cutoff frequency was manually optimized for each pulsar and arc based on the noise characteristics.
@@ III.4 Scintillation Arc Search and Curvature Measurement
[…]
Note. — The question mark in ”Arc ID F (?)” denotes that the grouping of these observations is tentative. For arc group F, due to limited observation data and significant curvature variations, we cannot confirm whether these arcs originate from the same screen.
Figure 2: Ten scintillation arcs were observed in pulsar B1237+25. Among them, nine scintillation arcs (excluding arc I) were identified during the observation on January 17, 2020. The left panel shows the secondary spectrum of that day, while the right panel shows the fitted scintillation arcs, labeled from A to J (excluding I) in order of increasing curvature.
@@ IV Results
We observed scintillation arcs in all observations (as shown in Table 3, 4, and 5). The observed scintillation arcs varied in structure, with some being more diffuse (e.g., PSR B1842+14, as shown in Figure 4) and others being finer (e.g., the scintillation arc G of PSR B1237+25, as shown in Figure 2 and 3). We note that except for arc F of pulsar B2021+51, all other scintillation arcs exhibited sharp parabolic arc characteristics, consistent with strongly anisotropic scintillation screens. For each anisotropic arc, we modeled its annual modulation using the Markov Chain Monte Carlo (MCMC) method to calculate the fractional distance $s$ of the scintillation screen, the velocity component of the scintillation screen along the scattering direction $v_{{\text{screen},~\zeta}}$, and the scattering direction $\mathbf{n}_{\text{screen}}$, as shown in Table 6. We express the measured direction of the scattering direction using the azimuth angle $\zeta$ as degrees east of north. The MCMC sampling was performed using the Python package emcee (Foreman-Mackey et al., 2013), which efficiently explores the parameter space to generate posterior distributions for the parameters.
@@
However, we were unable to calculate all three parameters for every scintillation screen when modeling the modulation. As discussed in Section II, in cases of anisotropic scattering, the direction of the scattering direction of the scintillation screen is the most easily constrained parameter, as it only requires knowledge of the point of minimum curvature. The fractional distance $s$ is the next most easily constrained, while the screen’s projected velocity $v_{{\text{screen},~\zeta}}$ is the most difficult to constrain. For some screens, we could only calculate partial parameters based on certain assumptions. Below are the situations we encountered and our approaches, respectively:
@@
1. Substantial errors. The relative errors in the curvature measurements of the scintillation arc H for PSR B1237+25 are substantial, resulting in poor constraints on the screen parameters. In this case, considering that the projected velocity $v_{{\text{screen},~\zeta}}$ of the screen is often relatively small and is the most difficult to constrain, we assumed its value to be 0 and, after modeling, could only obtain estimates of $s$ and $\zeta$.
@@
2. Weak modulation. When the scintillation screen is close to the pulsar, i.e., $s{\mathbf{v}_{\text{Earth}}}\cdot\mathbf{n}_{\text{screen}}\ll(1-s)\mathbf{v}_{\text{psr}}\cdot\mathbf{n}_{\text{screen}}$, the modulation amplitude of the arc curvature is not significantly greater than the measurement error. For example, the difference between the maximum and minimum curvature of scintillation arc A for PSR B1237+25 is only 0.15 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$, while the measurement error ranges from 0.05 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$ to 0.13 $\mathrm{m}^{-1}\,\mathrm{mHz}^{-2}$ (see Table 3). Scintillation arcs D and F of PSR B1237+25 (see Table 3) and scintillation arcs A, B, and C of PSR B2021+51 (see Table 5) also exhibit this characteristic. In such cases with small modulation amplitude, we could only constrain one screen parameter. Considering that $v_{{\text{screen},~\zeta}}$ is often relatively small, and these scintillation screens are relatively close to the pulsar, satisfying $v_{{\text{screen},~\zeta}}\ll(1-s)\mathbf{v}_{\text{psr}}\cdot\mathbf{n}_{\text{screen}}$, we assumed $v_{{\text{screen},~\zeta}}$ to be 0. We also assumed the main scattering direction $\mathbf{n}_{\text{screen}}$ is aligned with the effective velocity (see e.g. Sprenger et al. 2020),
$\mathbf{v}_{\mathrm{eff}}\equiv-\frac{1}{s}\mathbf{v}_{\mathrm{screen}}+\mathbf{v}_{\mathrm{Earth}}+\frac{1-s}{s}\mathbf{v}_{\mathrm{psr}}.$ (13)
@@
3. Insufficient number of detections. Some scintillation arcs could only be identified in observations conducted on dates with close day-of-year positions. Scintillation arcs B and E of PSR B1237+25 (see Table 3) were only observed on January 17, 2020, and January 14, 2023, which is effectively equivalent to having only a single observation. In such cases, we assumed $v_{{\text{screen},~\zeta}}$ to be 0 and $\mathbf{n}_{\text{screen}}$ to be aligned with $\mathbf{v}_{\text{eff}}$ to calculate the value of $s$. We obtained two solutions for $s$: the maximum value of $s$ if the scintillation screen is close to the pulsar, and the minimum value of $s$ if the scintillation screen is close to Earth.
4. Insufficient sampling of the modulation curve. For scintillation arc I of PSR B1237+25, two of the four observations were conducted on dates with close day-of-year positions (November 21, 2019, and November 16, 2022), lacking sufficient phase information about the annual modulation of curvature. This made constraining the screen parameters difficult. In this case, we assumed $v_{{\text{screen},~\zeta}}$ to be 0, and obtained estimates of $s$ and $\zeta$. For PSR B1237+25, the scintillation arc C was observed twice with about a 129-day interval in day-of-year positions, and their measured curvature values were very close, even less than the measurement error. Arc C could be a weakly modulated arc, or it might be a strongly modulated arc that happened to be observed on two dates with very similar curvature values. Therefore, we assumed $v_{{\text{screen},~\zeta}}$ to be 0 and $\mathbf{n}_{\text{screen}}$ to be aligned with $\mathbf{v}_{\text{eff}}$ to calculate the value of s. We obtained two solutions for s. The scintillation arc F of PSR B2021+51 had fewer observation data points and large differences in measured curvature. We could not determine whether this group of scintillation arcs originated from the same screen, so we did not fit it.
@@
Table 6: Parameters of scintillation screens observed in the directions of pulsars B1237+25, B1842+14, and B2021+51. The parameter $s$ represents the fractional distance of the scintillation screen, $d_{\text{screen}}$ represents the distance from the scintillation screen to the Earth, $\zeta$ represents the azimuth angle (degrees east of north) of the scattering direction, and $v_{{\text{screen},~\zeta}}$ represents the velocity component along the scattering direction. For screens without a $\zeta$ value, we assume their scattering direction and effective velocity are aligned. For screens without a $v_{{\text{screen},~\zeta}}$ value, we assume that their $v_{{\text{screen},~\zeta}}$ is zero.
Pulsar Screen ID $s$ $d_{\text{screen}}$ (pc) $\zeta$ () $v_{{\text{screen},~\zeta}}$ (km/s)
B1237+25 A $0.0435_{-0.0027}^{+0.0031}$ $824^{+61}_{-53}$ - -
@@
F $0.512^{+0.018}_{-0.017}$ $421^{+35}_{-31}$ - -
G $0.691^{+0.026}_{-0.028}$ $267^{+32}_{-28}$ $124.7^{+5.1}_{-5.2}$ $1.2^{+8.6}_{-9.5}$
H $0.632^{+0.096}_{-0.151}$ $321^{+130}_{-88}$ $171.1^{+10.0}_{-10.5}$ -
I $0.756^{+0.114}_{-0.224}$ $214^{+191}_{-102}$ $175.1^{+11.6}_{-12.7}$ -
J $0.723^{+0.069}_{-0.099}$ $241^{+86}_{-63}$ $63.7^{+8.4}_{-7.3}$ $41^{+22}_{-37}$
B1842+14 - $0.85^{+0.05}_{-0.10}$ $2.4^{+2.1}_{-1.2}$ $\times 10^{2}$ $122^{+11}_{-12}$ $19^{+11}_{-11}$
B2021+51 A $0.0368^{+0.0062}_{-0.0047}$ $1.93^{+0.31}_{-0.24}$ $\times 10^{3}$ - -
B $0.0567^{+0.0093}_{-0.0072}$ $1.88^{+0.31}_{-0.23}$ $\times 10^{3}$ - -
C $0.100^{+0.015}_{-0.012}$ $1.80^{+0.29}_{-0.22}$ $\times 10^{3}$ - -
@@
E $0.556^{+0.041}_{-0.038}$ $887^{+167}_{-132}$ $20.3^{+4.3}_{-4.3}$ $12.1^{+3.7}_{-3.8}$
Figure 6: Screen distance estimates for B1237+25 (top), B1842+14 (middle), and B2021+51 (bottom). The scintillation screen labeled “No assumption” indicates their positions were precisely calculated with all three parameters determined. The scintillation screen labeled “One assumption” indicates their positions were calculated based on one assumption that the velocity is zero. The scintillation screen labeled “Two assumptions” indicates their positions were calculated based on two assumptions: zero velocity and scattering direction aligned with the effective velocity. For pulsars with multiple scintillation screens, each screen has been marked with its ID. For PSR B1237+25, scintillation screens B, C, and E have two sets of distance solutions. For solutions close to the Sun, we label them as $\text{B}_{1}$, $\text{C}_{1}$, and $\text{E}_{1}$, while for solutions close to the pulsar, we label them as $\text{B}_{2}$, $\text{C}_{2}$, and $\text{E}_{2}$. The black dashed line represents the distance of the Local Bubble inner surface (Pelgrims et al., 2020; Pelgrims, 2022). The distance uncertainties of the screens include contributions from the pulsar distance uncertainty. For screens near the pulsar, this correlation causes their error bars to overlap with the pulsar distance uncertainties. The fractional screen distances $s$ are provided in Table 6.
@@
IV.1 PSR B1237+25
We identified 10 scintillation arcs in the direction of pulsar B1237+25, as shown in Table 3. Among them, 9 scintillation arcs could be identified in the observation data from January 17, 2020, as shown in Figure 2. Previously, three scintillation arcs were detected by Main et al. (2022) using the MeerKAT Thousand Pulsar Array at 856 – 1712 MHz, one scintillation arc was detected by Fadeev et al. (2018) using the GBT and Arecibo radio telescope at 316 – 332 MHz, who also estimated the screen distance, and only a diffuse blend of power in the secondary spectrum was observed by Wu et al. (2022) using LOFAR at 120 – 180 MHz.
Scintillation arc G had the highest brightness and was identified in all 10 observations. We measured and modeled the curvature modulation of arc G (as shown in Figure 9). We calculated that the fractional distance $s$ of scintillation screen G is $0.691^{+0.026}_{-0.028}$ (corresponding to a distance of $267^{+32}_{-28}$ pc from Earth), the scattering direction $\zeta$ is $124.7^{+5.1}_{-5.2}$ , and the velocity of the scintillation screen along the scattering direction $v_{{\text{screen},~\zeta}}$ is $1.2^{+8.6}_{-9.5}$ km/s. Fadeev et al. (2018) estimated the distance to the only scintillation screen they detected in this direction to be 0.23 $\pm$ 0.05 kpc, which is consistent with the distance to screen G. We also modeled and measured the curvature modulation of scintillation arcs J and I, successfully calculating all three parameters.
The measurement error of the curvature of scintillation arc H was relatively large, making it difficult to constrain the scintillation screen parameters, especially the $v_{{\text{screen},~\zeta}}$. Therefore, we assumed $v_{{\text{screen},~\zeta}}$ to be 0 and only calculated the fractional distance $s$ and the scattering direction. Scintillation arc I was identified in only 4 observations, with two occurring on dates in close proximity, preventing us from obtaining sufficient phase information about the curvature modulation. Hence, we also assumed $v_{{\text{screen},~\zeta}}$ to be 0.